OWG Meeting 21-Sep-2016

From GlueXWiki
Jump to: navigation, search

Location and Time

Room: CC F326

Time: 1:30pm-2:45pm


You can connect using BlueJeans Video conferencing (ID: 120 390 084). (Click "Expand" to the right for details -->):

(if problems, call phone in conference room: 757-269-6460)

  1. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 120390084.
  2. To join via a Web Browser, go to the page https://bluejeans.com/120390084.
  3. To join via phone, use one of the following numbers and the Conference ID: 120390084
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

Previous Meeting


  1. Announcements
  2. Collaboration Meeting
  3. DAQ
    • New firmware installation status
    • Busy testing with fADC125
  4. RAID to silo data flow
  5. AOT

Recharge Wednesday: - Not served this week due to Cyber Physical Systems Summit -

Offsite option


Attendees: David L. (chair), Dave A., Graham H., Sean D., Naomi J., Carl T., Sergey F., Beni Z., Simon T., Curtis M., Bryan M.


  • ROOT6 is now the default version for hdops. Some performance issues reported. No show stoppers yet.

Collaboration Meeting

  • Draft agenda is up
  • Online section moved to end of meeting to reflect current priorities of collaboration
  • Hovanes and Alex S. will be notified of their slots


  • Firmware testing still underway
  • Word on the street is that Alex S. has a working set of libraries in a private area but it is not yet installed in global area used by hdops
  • Street also believes all modules have had their firmware updated
  • Busy testing of fADC125 has not been done by Sergey since he is waiting on the DAQ system to become available
  • CODA 3.0.6 is ready, but has had only minimal testing due to the DAQ system being unavailable. Default is still CODA 3.0.5
    • Will do more CODA 3.0.6 testing when/if we have time prior to starting the FALL run. May need to do this during down times.


  • Online monitoring system checkout underway
  • Some minor tweaks required with change to ROOT 6
  • Overnight run showed memory leak of ~2GB/hr
    • Will not be an issue for 2hr runs as planned, but may be if left on much longer (e.g. 6 hr)
    • Exhausting memory on farm nodes can cause them to stall in a way that requires a power cycle to recover
    • Will check on setting a ulimit to kill the process before this happens

RAID to silo data flow

  • Work is ongoing but no significant changes to plan since last meeting